OpenStudy (anonymous):

What is the product of 3 x squared over the sum of x and 2 · the difference of 5 x squared and 20 all over 30 x cubed? Simplify completely

OpenStudy (anonymous):

\[\frac{ 3x^2 }{ x+2 } \times \frac{ 5x^2 - 20 }{ 30x^3 }\]

OpenStudy (mathstudent55):

Factor the numerator of the right fraction.

OpenStudy (anonymous):

uhhhh... what?

OpenStudy (mathstudent55):

Can you factor a common factor out of 5x^2 and 20?